OpenStudy (s):

How to sketch the graph of f(x)=x^2(x^2-4) including the inflection point?

OpenStudy (anonymous):

start with \[y=x^2(x-2)(x+2)\] so you see that you have 3 zeros, at x =-2,x=0,x=2

OpenStudy (anonymous):

the zero at x = 0 has multiplicity 2, so the graph will touch the x - axis there but not cross it

OpenStudy (anonymous):

to find the point of inflection you need the second derivative, \[f(x)=x^4-4x^2\] \[f'(x)=4x^3-8x\] \[f''(x)=12x^2-8\] set equal to zero and solve

OpenStudy (s):

My main question is that, when i find the inflection point I put square root of 2/3 into the original formula and i get -20/3 right?

OpenStudy (s):

And this point can not be on the graph, this makes me confused

OpenStudy (anonymous):

yes, for the second coordinate you evaluate the function at \[\sqrt{\frac{2}{3}}\]

OpenStudy (s):

yes, so I should get 2 inflection points

OpenStudy (anonymous):

it must be on the graph, plug in the x, find the y

OpenStudy (anonymous):

right \[\pm\sqrt{\frac{2}{3}}\]

OpenStudy (s):

so when I plug I get - 20/3

OpenStudy (anonymous):

for point of inflection we need only to equate first derivative of f(x) to 0.not the second

OpenStudy (anonymous):

actually it is the second first derivative gives relative max or min

OpenStudy (s):

it should be plugged in the original formula, right?

OpenStudy (anonymous):

yes, original function you have the x, you want the y

OpenStudy (s):

correct

OpenStudy (s):

But in the end my inflection point comes up to be (- square root of 2/3 ; - 20/3) and (square root of 2/3 ; - 20/3) and my minimum points are (-2;-4) and (2;-4) and it seems that the point of inflection cant be on the graph

OpenStudy (anonymous):

but im sure that i can get the points of inflection from first derivative,not the second